Explain the graph in the Solow Model (exogenous growth model) that relates capital intensity and output per worker.

What is the next period’s capital stock in the Solow Model? Explain it.
Answer #1

Solow Growth Model explains how growth in capital stock, in labor force and technology advancement interacts in an economy and how they affect the nation's output.

Solow Model is based on Neo Classical Theory of Economics which focusses on the Capital Accumulation process.

A change in the capital stock = s*f(k) - delta*k

Next period capital stock - previous period capital stock = s*f(k) - delta*k

So, Next period capital stock = s*f(k) - delta*k + previous period capital stock

So, the capital follows the process of capital accumulation given by neo- classical theory. Over time, the capital will accumulate and the previous stock of capital will get added to the new capital stock.